# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v2

@@ Abstract
Real-time prediction of corridor-level crash occurrence risk (COR) remains challenging, as existing near-miss based extreme value models oversimplify collision geometry, exclude vehicle–infrastructure (V–I) interactions, and inadequately capture spatial heterogeneity in vehicle dynamics. This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a nonstationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors. High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme near misses (V–V and V–I). The model incorporates dynamic variables and road features as covariates, with partial pooling across locations to address unobserved heterogeneity. The results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing the DIC by up to 7.5% for V–V and 3.1% for V–I near-misses. Predictive validation using ROC–AUC confirms strong performance: 0.89 for V–V segments, 0.82 for V–V intersections, 0.79 for V–I segments, and 0.75 for V–I intersections. Model interpretation reveals that relative speed and distance dominate V–V risks at intersections and segments, with deceleration critical in segments, while V–I risks are driven by speed, boundary proximity, and steering/heading adjustments. These findings highlight the value of a statistically rigorous, geometry-sensitive, and spatially adaptive modeling approach for proactive corridor-level safety management, supporting real-time interventions and long-term design strategies aligned with Vision Zero.
@@ 1 Introduction
The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) require innovative safety assessment approaches. Urban corridors are particularly vulnerable, where a single incident can disrupt mobility, trigger cascading congestion, and impose severe economic losses. In 2023, traffic crashes in the United States claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes is estimated to exceed \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite extensive research on crash patterns, risk estimation, and mitigation strategies (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b), accurate prediction of crash risk in urban corridors remains methodologically underdeveloped.
Most existing models focus exclusively on vehicle–vehicle (V–V) near-misses, largely due to the availability of vehicle trajectory data, while systematically excluding vehicle–infrastructure (V–I) interactions such as lane edges, curbs, barriers, and crosswalks. However, V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021). This omission makes corridor-level safety models inherently incomplete. Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now allow the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
Crash risk prediction strategies based on historical retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011) suffer from well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time conditions. These challenges motivate a paradigm shift towards proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). Using high-resolution vehicle trajectory data, SSMs reflect the underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among SSMs, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, these formulations often assume constant velocity and linear trajectories, leading to biased results in boundary-constrained environments. Recent research has extended TTC to two-dimensional (2D) indicators of higher-fidelity (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration. However, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence the likelihood and severity of near misses. To support a reliable short-term crash occurrence risk (COR) estimate, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure. This study builds on these efforts by introducing a high-fidelity, geometry-aware 2D-TTC indicator that explicitly incorporates vehicle geometry, dynamic states, and roadway boundaries, ensuring realistic and efficient detection of both V–V and V–I near misses.
@@
While SSMs capture frequent near-misses, a robust COR estimation framework requires extrapolating extreme near-misses to rare event outcomes. To date, several studies have proposed probabilistic (Saunier and Sayed, 2008), causal (Davis et al., 2011), and extreme value theory (EVT) (Fu and Sayed, 2021a, b; Zheng et al., 2021b; Ali et al., 2023a; Zheng et al., 2021a; Songchitruksa and Tarko, 2006; Zheng et al., 2014; Tarko, 2012). EVT provides a principled statistical foundation for this task, modeling the tail behavior of the extreme near-miss distribution. Two EVT approaches dominate: the Generalized Extreme Value (GEV) distribution using Block Maxima (BM) sampling and the Generalized Pareto Distribution (GPD) using Peaks over Threshold (POT) sampling (Coles et al., 2001). The BM approach is particularly suitable for real-time analysis due to its short block structure, lower threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
@@
Early applications of EVT in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area. More recently, non-stationary GEV models have been introduced, allowing covariates to influence location and scale parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby capture heterogeneity across traffic states (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions have also addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). However, most applications remain geographically narrow and functionally limited to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024). Although some studies have incorporated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), these frameworks generally omit explicit integration of roadway geometry and directional heterogeneity at the corridor level. Such integration requires a modeling structure capable of jointly representing vehicle dynamics (which vary temporally) and roadway features (which vary spatially across segments and intersections). Therefore, a critical gap remains for a comprehensive Bayesian hierarchical GEV framework capable of jointly modeling vehicle dynamics, roadway geometry, flow directionality, and multilevel heterogeneity. This framework is crucial for accurately estimating the COR at the corridor level, which reflects the complex spatiotemporal dynamics of urban traffic environments. This paper makes the following unique contributions:
  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in previous studies.
  • Introduces a geometry-aware high-fidelity 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
@@
  • Advances a hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) Univariate GEV model that integrates roadway geometry, traffic flow directionality, and multilevel heterogeneity, enabling corridor-wide yet location-sensitive crash risk estimation
@@
Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for a proactive COR framework in urban corridors, leveraging Arogovers-2 AV trajectory and HD map data to support both operational interventions and long-term safety planning.
@@ 4 Data description
To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and D.C.) using SAE Level 4 AVs. Argoverse-2 comprises more than 250k driving scenarios that span 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions, such as intersection near-misses, lane changes, and pedestrian crossings, conditions most indicative of crash risk. A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic. Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors. The sensor suite includes seven ring cameras and two front-facing stereo cameras (20 Hz) supported by dual 64-beam VLP-32C LiDAR sensors, producing  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
@@ 4.1.1 Motion forecasting dataset
Each motion forecasting scenario follows a hierarchical structure with three levels: scenario metadata, agent tracks, and frame-level states. Scenarios are 11s duration and are sampled at 10 Hz, producing 110 time steps per vehicle for fine-grained interaction modeling. At the scenario level, metadata include a unique scenario ID, the recording city, and the focal agent, typically the ego AV. Each scenario then contains multiple agent tracks, where each track corresponds to a road user and specifies the object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal, scored, or unscored), and a time-ordered sequence of states. Frame-level states capture position, velocity, and heading at each timestep, forming the basis for computing 2D-TTC.
@@
Because positions were reported in local, city-specific frames, all coordinates were transformed to globally referenced WGS84. This involved (i) projecting each city’s local origin into its UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), and (ii) offsetting local $(x,y)$ positions and inverse-projecting them to latitude–longitude using pyproj. This geospatial normalization provided consistent referencing across cities and ensured compatibility with HD map overlays. Batch scripts automatically traversed scenario folders, performed attribute extraction and coordinate conversion, and exported unified CSV datasets containing both original city-frame and converted lat/lon values. These datasets were then used for spatial filtering, interaction detection, and overlay validation.
@@ 4.2 Study area
Lane configurations vary between 2 and 3 lanes per direction, with average widths ranging from 3.1m (S7) to 3.6m (S6), indicating subtle but important differences in cross-sectional design. Driveway and minor intersection density also differs across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses. Median type alternates between raised (e.g., S1–S2) and flush paved (e.g., S3–S9), influencing turning maneuvers and midblock encroachments. Overall, the table highlights that intersections serve as near-miss-prone nodes with concentrated but lower-volume interactions, whereas segments provide richer samples of continuous traffic flow, characterized by greater geometric variability.
@@
After identifying and compiling the BM near-miss events for each observation unit (segments and intersections), the frequencies of extreme V–V and V–I events were systematically aggregated and visualized in Fig. 4.2. This plot illustrates the spatial distribution of extreme near-miss frequencies across 28 locations, comprising 10 intersections and 18 segments, organized into intersection zones, primary-direction segments, and opposite-direction segments. This spatial aggregation confirms the integrity of the BM dataset and provides the foundation for modeling. The study extends prior approaches by incorporating both V–V and V–I near misses into the HBSGRP–UGEV framework, explicitly integrating roadway geometry and infrastructure features alongside vehicle dynamics. Prior research has emphasized the role of vehicle kinematics in crash risk (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015); here, additional geometric covariates, such as lane configuration and boundary proximity, are introduced. For example, high steering angles, often indicating abrupt lateral maneuvers—were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021). Fig. 4.2 highlights several notable spatial patterns. Opposite-direction segments (S82, S72, S62) exhibited the highest V–I frequencies, suggesting greater lateral variability and exposure to curb or median near-misses. Primary-direction segments (S11, S21, S81) showed elevated V–V near-miss frequencies, consistent with midblock environments dominated by longitudinal interactions such as close-following, lane changes, or overtaking. Among intersections, I7, I8, and I9 emerged as critical hotspots, with elevated frequencies of both near-miss types. This dual concentration reflects the operational complexity of intersections, where near-missing maneuvers, pedestrian flows, and signal control interact to amplify crash potential.

### v3

@@ Abstract
Real-time corridor-wide crash occurrence risk (COR) prediction is challenging because existing near-miss EVT models oversimplify collision geometry, neglect vehicle–infrastructure (V–I) interactions, and fail to capture spatial heterogeneity in traffic and roadway conditions. To address these gaps, this study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator for capturing V–V and V–I near misses, and a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) UGEV model for estimating short-term COR in urban corridors. High-resolution trajectories from the Argoverse-2 dataset were analyzed across 28 sites on Miami’s Biscayne Boulevard to extract extreme near-miss events. The model incorporates vehicle dynamics and roadway features as covariates, with partial pooling across segments and intersections to capture corridor-wide heterogeneity. Results show that the HBSGRP–UGEV framework outperforms fixed-parameter models, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near misses. Predictive validation using ROC–AUC confirms strong accuracy (0.89 for V–V segments, 0.82 for intersections, 0.79 for V–I segments, and 0.75 for intersections). Random-parameter effects showed that relative speed, distance, and deceleration dominated V–V near misses at segments (only distance for V–I), while intersections were dominated by relative speed and distance for V–V, with no significant effects for V–I.. These findings demonstrate the value of a geometry-aware, spatially adaptive framework for proactive corridor safety management, supporting both real-time interventions and long-term Vision Zero goals.
@@ 1 Introduction
The increasing complexity of urban corridors, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches. Urban corridors are particularly vulnerable, as even a single crash can disrupt mobility, trigger cascading congestion, and result in substantial economic losses. In 2023, traffic crashes in the United States claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes exceeds \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite decades of research, corridor-wide crash risk prediction remains methodologically limited.
Crash risk prediction has traditionally relied on retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011). While such approaches have advanced understanding of crash patterns, they face well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time traffic conditions. These drawbacks highlight the need for proactive approaches that anticipate rather than react to crash risk.
To address these challenges, researchers have shifted toward proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). SSMs use frequent near-miss events as proxies for crashes, enabling the study of traffic conflict dynamics at a much higher temporal resolution (Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among these, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, their assumptions of constant velocity and linear paths bias results in boundary-constrained settings. Recent two-dimensional (2D) extensions (Li et al., 2024; Anis et al., 2025b) incorporate position, heading, and acceleration, yet still rely on simplified circular vehicle approximations. This study addresses these limitations by introducing a geometry-aware 2D-TTC that accounts for vehicle dimensions, dynamics, and roadway boundaries to realistically capture both vehicle–vehicle (V–V) and vehicle–infrastructure (V–I) near misses.
Building on this foundation, recent studies have integrated SSMs with extreme value theory (EVT) to estimate crash occurrence risk (COR) in urban corridors (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). EVT extrapolates from frequent near misses to rare crash-prone extremes, offering a statistically principled framework for proactive safety analysis. Yet most SSM–EVT applications remain limited, focusing narrowly on V–V interactions and excluding V–I conflicts such as lane edges, curbs, and barriers. This omission is critical because V–I events account for a large share of severe roadway departure crashes (Islam et al., 2021). Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable precise alignment of roadway boundaries with vehicle trajectories, creating new opportunities to model both V–V and V–I near misses with geometric fidelity.
@@
While SSMs capture frequent near misses, robust COR estimation requires extrapolating extreme near misses to rare crash outcomes. EVT (Fu and Sayed, 2021a, b; Zheng et al., 2021b; Ali et al., 2023a; Zheng et al., 2021a; Songchitruksa and Tarko, 2006; Zheng et al., 2014; Tarko, 2012) has become the most widely used method for this task, with two dominant formulations: the Generalized Extreme Value (GEV) distribution using Block Maxima (BM) sampling, and the Generalized Pareto Distribution (GPD) using Peaks-over-Threshold (POT) sampling (Coles et al., 2001). The BM approach is particularly well-suited to real-time applications due to its short block structure, reduced threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
@@
Early applications of EVT in traffic safety used stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022), which assumed constant parameters across study areas. More recent work has developed nonstationary GEV formulations, allowing covariates to influence distribution parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and capture heterogeneity across traffic states. (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions further address data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). Despite these advances, most applications remain geographically narrow and functionally restricted to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024). Although some studies have integrated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), explicit corridor-wide models that jointly integrate vehicle dynamics, roadway geometry, and directional heterogeneity remain rare.
@@
To address this gap, this paper makes three contributions: (1) it develops a corridor-wide COR framework that incorporates both V–V and V–I interactions, (2) it introduces a geometry-aware 2D-TTC indicator that captures extreme near-miss dynamics with higher geometric fidelity, and (3) it advances a hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) GEV model that integrates roadway geometry, traffic directionality, and multilevel heterogeneity for corridor-wide yet location-sensitive crash risk estimation.
@@
Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for proactive corridor-wide crash risk prediction, leveraging Argoverse-2 trajectory and HD map data to support both real-time interventions and long-term safety planning.
@@ 4 Data description
To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and D.C.) using SAE Level 4 AVs. Argoverse-2 comprises more than 250k driving scenarios that span 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions, such as intersection near-misses, lane changes, and pedestrian crossings, conditions most indicative of crash risk. A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic. Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors. The sensor suite comprises seven ring cameras and two front-facing stereo cameras (20 Hz), supported by dual 64-beam VLP-32C LiDAR sensors, which produce  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
@@ 4.1.1 Motion forecasting dataset
Each motion forecasting scenario follows a hierarchical structure with three levels: scenario metadata, agent tracks, and frame-level states. Scenarios are 11s duration and are sampled at 10 Hz, producing 110 time steps per vehicle for fine-grained interaction modeling. At the scenario level, metadata includes a unique scenario ID, the recording city, and the focal agent, typically the ego AV. Each scenario then contains multiple agent tracks, where each track corresponds to a road user and specifies the object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal, scored, or unscored), and a time-ordered sequence of states. Frame-level states capture position, velocity, and heading at each timestep, forming the basis for computing 2D-TTC.
@@
Because positions were reported in local, city-specific frames, all coordinates were transformed into a globally referenced WGS84 system. This involved (i) projecting each city’s local origin into its UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), and (ii) offsetting local $(x,y)$ positions and inverse-projecting them to latitude–longitude using pyproj. This geospatial normalization provided consistent referencing across cities and ensured compatibility with HD map overlays. Batch scripts automatically traversed scenario folders, performed attribute extraction and coordinate conversion, and exported unified CSV datasets containing both original city-frame and converted lat/lon values. These datasets were then used for spatial filtering, interaction detection, and overlay validation.
@@ 4.2 Study area
Lane configurations vary between 2 and 3 lanes per direction, with average widths ranging from 3.1m (S7) to 3.6m (S6), indicating subtle but important differences in cross-sectional design. Driveway and minor intersection density also differ across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses. The median type alternates between raised (e.g., S1–S2) and flush-paved (e.g., S3–S9), influencing turning maneuvers and mid-block encroachments. Overall, the table highlights that intersections serve as near-miss-prone nodes with concentrated but lower-volume interactions, whereas segments provide richer samples of continuous traffic flow, characterized by greater geometric variability.
@@
After identifying and compiling the BM near-miss events for each observation unit (segments and intersections), the frequencies of extreme V–V and V–I events were systematically aggregated and visualized in Fig. 4.2. This plot illustrates the spatial distribution of extreme near-miss frequencies across 28 locations, comprising 10 intersections and 18 segments, organized into intersection zones, primary-direction segments, and opposite-direction segments. This spatial aggregation confirms the integrity of the BM dataset and provides the foundation for modeling. The study extends prior approaches by incorporating both V–V and V–I near misses into the HBSGRP–UGEV framework, explicitly integrating roadway geometry and infrastructure features alongside vehicle dynamics. Prior research has emphasized the role of vehicle kinematics in crash risk (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015); here, additional geometric covariates, such as lane configuration and boundary proximity, are introduced. For example, high steering angles, often indicating abrupt lateral maneuvers, were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021). Fig. 4.2 highlights several notable spatial patterns. Opposite-direction segments (S82, S72, S62) exhibited the highest V–I frequencies, suggesting greater lateral variability and exposure to curb or median near-misses. Primary-direction segments (S11, S21, S81) showed elevated V–V near-miss frequencies, consistent with midblock environments dominated by longitudinal interactions such as close-following, lane changes, or overtaking. Among intersections, I7, I8, and I9 emerged as critical hotspots, with elevated frequencies of both near-miss types. This dual concentration reflects the operational complexity of intersections, where near-missing maneuvers, pedestrian flows, and signal control interact to amplify crash potential.